OpenStudy (anonymous):

Find the intercepts 5y-15=5x

OpenStudy (anonymous):

would you be willing to set y to be zero and solve for x?

OpenStudy (anonymous):

that will give your the x intercept

OpenStudy (anonymous):

then, to get the y intercept set the x to zero and solve for y

OpenStudy (anonymous):

Simplifying 5y + -15 = 5x Reorder the terms: -15 + 5y = 5x Solving -15 + 5y = 5x Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Add '15' to each side of the equation. -15 + 15 + 5y = 15 + 5x Combine like terms: -15 + 15 = 0 0 + 5y = 15 + 5x 5y = 15 + 5x Divide each side by '5'. y = 3 + x Simplifying y = 3 + x

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

you make me look really lazy lol

OpenStudy (anonymous):

sorry?

OpenStudy (anonymous):

that is a nice explanation is what i meant

OpenStudy (anonymous):

opps my mistake, then you can graph!

OpenStudy (anonymous):

oh haha

OpenStudy (anonymous):

lol thanks you two :D

OpenStudy (anonymous):

np :D need help with graphing?

OpenStudy (anonymous):

no, do I do the same to find the x?